1. if you have 3 / 4 cup of flour how many eighths do you have?
2. if you 3/12 of an orange how many fourths do you have?

Answer by Theo(13342)

set up a ratio.
the first ratio is:
3/4 = x/8
cross multiply to get 4x = 24
divide both sides by 4 to get x = 6
your ratio becomes 3/4 = 6/8
the second ratio is:
3/12 = x/4
cross multiply to get 12x = 12
divide both sides by 12 to get x = 1
your ratio becomes 3/12 = 1/4

another way of solving this is as follows:
for the first problem:
you want to know how many eighths in 3/4.
since 8 is 2 times 4, multiply numerator and denominator by 2 to get 3/4 = 6/8
this is because 2 * 3 = 6 (numerator) and 4 * 2 = 8 (denominator)
for the second problem:
you want to know how many fourths in 3/12.
since 12/3 = 4, you want to divide both numerator and denominator by 3 to get 3/12 = 1/4.
this is because 3 / 3 = 1 (numerator) and 12 / 3 = 4 (denominator).

in both cases, you want to know what you want to do to the denominator to make the new denominator the value requested.

in the first problem your original denominator is 4 and you have to multiply it by 2 to make the new denominator equal to 8.

in the second problem your original denominator is 12 anbd you have to divide it by 3 to make the new de3nominator equal to 4.

in order to keep the fraction the same value, you have to multiply or divide the numerator by the same factor.

if you multiply the denominator by 2, you have to also multiply the numerator by 2 to keep the fraction the same.

if you divide the denominator by 4, you have to also divide the numerator by 4 to keep the fraction the same.